# Dynamical cluster-based optimization of tensor network algorithms for quantum circuit simulations

### v1

@@ Abstract
We optimize Matrix-Product State (MPS)-based algorithms for simulating quantum circuits with finite fidelity, specifically the Time-Evolving Block Decimation (TEBD) and the Density-Matrix Renormalization Group (DMRG) algorithms, by exploiting the irregular arrangement of entangling operations in circuits. We introduce a variation of the standard TEBD algorithm, we termed “cluster-TEBD”, which dynamically arranges qubits into entanglement clusters, enabling the exact contraction of multiple circuit layers in a single time step. Moreover, we enhance the DMRG algorithm by introducing an adaptive protocol which analyzes the entanglement distribution within each circuit section to be contracted, dynamically adjusting the qubit grouping at each iteration. We analyze the performances of these enhanced algorithms in simulating both stabilizer and non-stabilizer random circuits, with up to $1000$ qubits and $100$ layers of Clifford and non-Clifford gates, and in simulating Shor’s quantum algorithm with tens of thousands of layers. Our findings show that, even with reasonable computational resources per task, cluster-based approaches can significantly speed up simulations of large-sized quantum circuits and improve the fidelity of the final states.
@@ I Introduction
Quantum computing promises significant performance advantages over classical algorithms, particularly for tasks such as search [1] and factorization [2]. However, quantum computers still face serious challenges, as quantum states are vulnerable to physical effects, such as noise and decoherence, that lead to loss of information over time [3, 4]. Although significant progress has been made recently in designing efficient error-correction codes [5], accurately executing long-time evolutions on quantum computers remains a significant challenge.
@@
At the same time, simulating a quantum system on a classical computer poses its own challenges. In fact, a generic quantum state of a qubit system contains an amount of information that grows exponentially with the number of qubits. Consequently, the memory required to exactly represent such quantum states quickly exceeds the capabilities of any existing classical computer. This problem can be mitigated through the use of tensor network techniques [6, 7, 8]. In one-dimensional systems, quantum states can be efficiently represented as Matrix Product States (MPS) [9, 10], which require significantly less memory than a full quantum state description, albeit at the cost of introducing approximation errors. By leveraging efficient MPS-based algorithms, such as Time-Evolving Block Decimation (TEBD) [11] and the Density Matrix Renormalization Group (DMRG) [12, 13], it is possible to achieve very low approximation errors and numerically address key problems in the study of static and dynamical properties of quantum many-body systems.
@@
The utility of tensor networks in quantum many-body physics has rapidly extended to quantum computing. While simulating ideal large-scale quantum computers on classical machines is exceedingly challenging, noisy intermediate-scale quantum (NISQ) computers exhibit imperfections that lead to computational errors comparable with the approximation errors of tensor network algorithms [14]. Following Google’s claim of achieving quantum supremacy [15] by simulating a random quantum circuit [16, 17, 18, 15] with a superconducting quantum device [19, 20], a series of studies demonstrated that it was possible to simulate the same task with similar or better fidelity on large classical computers with optimized tensor network contraction techniques [21, 22, 23, 24]. Another work by the team of IBM Quantum showed evidence for the utility of NISQ devices [4] by simulating the 2D transverse-field Ising model with a Trotterized quantum circuit on a 127-qubit quantum computer [25]. Shortly thereafter, tensor network simulations of the same Hamiltonian were performed with good accuracy on relatively low computational resources, increasing the simulated amount of qubits up to 1121 [26, 27].
@@
Optimized classical simulations of many-body dynamics and quantum computing can be seen as complementary in the study of collective quantum effects. Indeed, optimizing numerical methods to simulate large and complex quantum systems could lead to understanding quantum phenomena at large scales, thereby aiding in the development of quantum technologies, such as creating digital twins of experimental quantum platforms [28, 29], or optimizing TEBD and DMRG algorithms to simulate both random quantum circuits and quantum algorithms [14, 30, 31].
@@
In this work, we propose an approach based on entanglement clustering [32, 33, 34] to improve tensor network algorithms for simulating quantum circuits with finite fidelity. The main idea is that, by dynamically clustering qubits, according to the pattern of entangling gates in a quantum circuit, we can optimize the simulation of deep quantum circuits with irregular entanglement distribution.
@@
In particular, we focus on the TEBD and DMRG algorithms. For the former, instead of immediately applying individual gates to the initial MPS, we perform an analysis of the structure of the simulated quantum circuit as a whole. We adaptively divide a quantum circuit into clusters of entanglement according to the amount of memory available in our classical computing system, allowing the algorithm to exactly contract multiple gate layers at the same time. We call this variant the “cluster-TEBD” algorithm. For the latter, we devise a protocol that makes use of hypergraph partitioning [35] to automatically determine how to group qubits throughout the algorithm, and we exploit this protocol to adaptively vary the grouping in each iteration, based on the distribution of entangling gates in the contracted section of the circuit. Moreover, we employ heuristic approaches to contract multiple layers efficiently [36, 21].
@@
Benchmarking the original algorithms against their respective optimized variations on random circuits, with both Clifford and non-Clifford gates [37], as well as on a tensor network-based implementation of Shor’s algorithm for factoring [2], shows substantial improvements in both runtime and fidelity, proving the effectiveness of cluster-based optimization in finite-fidelity tensor network algorithms.
@@
The paper is organized as follows. In Sec. II, we make a comprehensive review of tensor network methods to simulate quantum circuits with finite fidelity. In Sec. III, we explain how arranging qubits in clusters of highly-entangled subsystems can improve TEBD and DMRG simulations of quantum circuits. In Sec. IV, we present the results of our benchmarks, where we compare runtime and fidelity of cluster-TEBD and standard TEBD when simulating random-structure quantum circuits and Shor’s algorithm, and show the effect of contracting more layers in a single iteration of DMRG. Finally, in Sec. V, we draw our conclusions and discuss possible future directions.
@@ II Simulating quantum circuits with tensor networks
Figure 1: Quantum circuit with $8$ qubits and $7$ layers modelled as a tensor network. The initial state is an outer product of rank-1 tensors (vectors), as in this case it is a fully separable state. The gates are modelled as rank-2 tensors (matrices) for single-qubit operations, and rank-4 tensors for two-qubit operations, and they are labelled according to the qubits they act on and the layer they belong to.
In our computational framework, a tensor can be interpreted as a multidimensional array, with a number $r$ of indices, called the “rank” of the tensor, each one taking a finite number of values (“dimension”). Contracting tensors over their shared indices generates what is referred to as a “tensor network” [6, 7]. The purpose of this section is to introduce notation and briefly illustrate how tensor networks can be used to represent and simulate quantum circuits. In particular, we present some recent developments in the context of tensor network algorithms to simulate quantum circuits [14, 30], introducing some quantities that will be crucial throughout the text.
@@ II.1 Mapping quantum circuits to tensor networks
$N$-qubit gates in a quantum circuit can be directly represented as rank-$2N$ tensors with all indices of dimension 2. This means, for instance, that a two-qubit gate, which is usually represented as a $4\times 4$ matrix, can be mapped to a $2\times 2\times 2\times 2$ tensor. When no qubits are entangled at the start, the initial state of a circuit can also be directly represented as an outer product of rank-1 tensors (“vectors”), each of physical dimension $d=2$, corresponding to the dimension of the Hilbert space of a qubit.
@@
When dealing with high-depth quantum circuits, it is crucial to define a quantity that characterizes the position of each gate within the overall evolution. For this reason, we define a gate layer $l$ in a quantum circuit as the subsets of gates that operate on independent sets of qubits at most once at the same time step. Let us identify as $U_{(q_{1},\dots,q_{n}),l}$ a generic $n$-qubit gate applied on qubits $q_{1},\dots,q_{n}$ at layer $l$. For instance, the gates $U_{(3,4),l}$ and $U_{4,l^{\prime}}$ cannot belong to the same layer, as they operate on the same qubit $4$, making $l\neq l^{\prime}$. On the other hand, gates $U_{(3,4),l}$ and $U_{5,l^{\prime\prime}}$ could belong to the same layer, provided no operations are on qubit $5$ after $U_{(3,4),l}$. If this condition is satisfied, we can write $l=l^{\prime\prime}$. Gates belonging to the same layer can be physically interpreted as unitary operations applied at the same time step. Fig. 1 shows an example of a quantum circuit modelled as a tensor network, using the described notation for layers. Because of the way a gate layer was defined, the number of gates per layer can change depending on the structure of the circuit. In the circuit shown in Fig. 1, all layers are fully filled with gates, but if, for instance, we added two gates $U_{(6,7)}$ and $U_{(7,8)}$ at the end of the evolution, each of them would belong to two new separate layers $8$ and $9$, respectively.
Figure 2: Entangled $N$-qubit quantum state |$\ket{\psi}$, initially represented as a rank-$N$ tensor with total size $2^{N}$, decomposed into an $N$-site matrix-product state with successive singular value decompositions.
II.2 Encoding entanglement: bond dimensions, truncation and fidelity
@@
When the initial state of a quantum circuit is entangled, mapping it to tensor network notation is not as straightforward. Since the individual sites are not fully separable into tensor products, the entire state has to be stored in a tensor of size $2^{N}$, with $N$ being the number of qubits. However, such a tensor can be decomposed, as shown in Fig. 2, by performing successive singular value decompositions (SVD) to obtain a matrix-product state (MPS) representation [9, 10]. Aside from the physical dimensions, this representation introduces additional virtual bond dimensions $\chi_{b}$, which encode the entanglement of the state. Assuming for simplicity $\chi_{1}=\chi_{2}=\dots=\chi$, such a representation, in principle, contains a number of parameters equal to $2N\chi^{2}$, i.e., from exponential to linear in the number of qubits. Note that a separable state can be represented equivalently as an MPS with bond dimensions $\chi_{b}=1\ \forall\ b$.
When considering the bipartition of a quantum system related to a specific bond $b$, the entanglement between the two subsystems $A$ and $B$ can be quantified via the von Neumann entanglement entropy $S(\rho_{A}):=-\text{Tr}\left(\rho_{A}\log\rho_{A}\right)$, where $\rho_{A}:=\text{Tr}_{B}\left(\lvert\psi\rangle\langle\psi\rvert\right)$ is the reduced density matrix of subsystem $A$. We can equivalently denote the entanglement entropy as $S_{b}$, in relation to the bond $b$ where the bipartition originates from. It can be proven that the bond dimension $\chi_{b}$ of an MPS at bond $b$ scales exponentially with the bipartite entanglement entropy in the corresponding bond [38, 39]:
@@
In general, unitary evolutions, as those performed in quantum circuits, can lead to volume-law scaling, where the entanglement entropy scales linearly with the system size, e.g. for a one-dimensional system of size $L$:
@@
In this case, the bond dimension scales exponentially with the system size. Inevitably, the longer the evolution, the larger the bond dimension, making exact simulations extremely challenging. An approximation can be introduced by reducing the bond dimension via the truncation of singular values in the SVD step, as in Fig. 3. The singular values to truncate at each step can be determined using two parameters, namely the cutoff $\eta$ and the maximum bond dimension $\chi_{\mathrm{max}}$. The cutoff $\eta$ determines the minimum magnitude a singular value must have to be kept, meaning all singular values smaller than $\eta$ are discarded. The maximum bond dimension $\chi_{\mathrm{max}}$ is the limit on the number of singular values to be kept, meaning if the bond dimension is larger than $\chi_{\mathrm{max}}$, all values in excess are discarded.
@@
$\mathcal{F}(n)=\lvert\braket{\psi^{\mathrm{MPS}}_{\mathrm{exact}}(n)}{\psi^{\mathrm{MPS}}_{\mathrm{trunc}}(n)}\rvert^{2}.$ (3)
However, knowing the exact representation of a highly entangled quantum state requires large bond dimensions, leading back to the original problem. A way to circumvent this problem is to approximate the fidelity using singular values, as shown in  [14]. Let us define the two-qubit fidelity $f_{i}$ as the ratio between the sum of singular values after and before truncation:
@@
Together with runtime, the total fidelity can be used as a benchmark to evaluate the efficiency and the accuracy of classical algorithms for simulating quantum dynamics [14].
@@ II.3 Algorithms for simulating quantum circuits
The main algorithms for simulating quantum circuits, originally developed for simulating static and dynamical evolution of quantum many-body systems, exploit the approximate fidelity introduced in Eq. (5) to measure the simulation error on large/highly entangled quantum states [14, 30].
@@
The Time-Evolving Block Decimation (TEBD) algorithm iterates over each gate in the evolution and applies it to the corresponding qubits. Each iteration of the TEBD, analyzing a generic gate $U^{[i,i+1,\dots]}_{l+1}$, can be divided into three main steps:
@@
    Orthogonalization: a sequence of QR decompositions is applied to the MPS tensors $A^{[j]}_{l}$ with $j~{}\neq~{}i$, bringing the state into a canonical form where all tensors to the left (right) of site $i$ satisfy left (right) isometry conditions. In this way, the norm of the state |$\ket{\psi^{\mathrm{MPS}}}$ represented by the MPS can be computed as $\braket{\psi^{\mathrm{MPS}}}{\psi^{\mathrm{MPS}}}~{}=~{}\text{Tr}\left(A^{[i]}_{l}{A^{[i]\dagger}_{l}}\right)$. We call site $i$ the “orthogonality center” of the MPS.
@@
A different approach that requires the contraction of multiple layers of the circuit is the Density Matrix Renormalization Group (DMRG) algorithm for quantum circuit simulations, introduced by Ayral et al. [30]. The algorithm starts by defining a number of layers $L_{\mathrm{max}}$ to optimize in each step at the same time. Then, the initial MPS with tensors $A^{[1]}_{0},\dots,A^{[N]}_{0}$ is mapped to a grouped MPS with tensors $\Psi^{[1]}_{0},\dots,\Psi^{[g]}_{0}$, with $g\leq N$ the number of grouped sites. In a grouped MPS, each site $\tau$ has multiple physical indices $d_{\tau}$, depending on the number of qubits mapped to that site. Each grouped site is a tensor that represents a higher-dimensional system with a Hilbert space of dimension $2^{d_{\tau}}$, preserving all the entanglement within this group exactly without any truncation. A grouped site can be formed by contracting the sites from the original MPS that are to be combined, according to a predefined grouping scheme. The grouping scheme is predetermined prior to the algorithm’s execution and kept fixed throughout the simulation, and it is based on considerations about the structure of the circuits to simulate and on the amount of available memory. For instance, qubits on which many entangling operations are applied should belong to the same group, but the size of each grouped site should not exceed a certain threshold, as the system would run out of memory. In each step $t$, the algorithm generates a random grouped MPS $M^{[1]{\dagger}}_{tL_{\mathrm{max}}},\dots,M^{[g]{\dagger}}_{tL_{\mathrm{max}}}$ with bond dimensions $\tilde{\chi}_{i}$. Let $\Psi^{[1]}_{tL_{\mathrm{max}}},\dots,\Psi^{[g]}_{tL_{\mathrm{max}}}$ denote the MPS of the evolved state after applying $tL_{\mathrm{max}}$ layers of the circuit. The objective of the algorithm is to variationally minimize the distance between the random grouped MPS and the exact MPS representation of the state after $tL_{\mathrm{max}}$ layers. For illustrative purposes, we describe the iteration of the algorithm at step $1$ in the following:
@@
  • Compute $f_{\tau}=\text{Tr}\left(F^{[\tau]}_{L_{\mathrm{max}}}F^{[\tau]{\dagger}}_{L_{\mathrm{max}}}\right)$ and replace $M^{[\tau]}_{L_{\mathrm{max}}}=\dfrac{F^{[\tau]}_{L_{\mathrm{max}}}}{f_{\tau}}$.
This iteration can be generalized for any step $t$, where the evolution from layer $(t-1)L_{\mathrm{max}}+1$ to layer $tL_{\mathrm{max}}$ is optimized (for further details, see Ref. [30]).

### v2

# Dynamical cluster-based strategy for improving tensor network algorithms in quantum circuit simulations

@@ Abstract
We optimize matrix-product state-based algorithms for simulating quantum circuits with finite fidelity, specifically the time-evolving block decimation (TEBD) and the density-matrix renormalization group (DMRG) algorithms, by exploiting the irregular arrangement of entangling operations in circuits. We introduce a variation of the standard TEBD algorithm, we termed “cluster-TEBD”, which dynamically arranges qubits into entanglement clusters, enabling the exact contraction of multiple circuit layers in a single time step. Moreover, we enhance the DMRG algorithm by introducing an adaptive protocol, which analyzes the entanglement distribution within each circuit section to be contracted, dynamically adjusting the qubit grouping at each iteration. We analyze the performances of these enhanced algorithms in simulating both stabilizer and nonstabilizer random-structured quantum circuits, with up to $1000$ qubits and $100$ layers of Clifford and non-Clifford gates, and in simulating Shor’s quantum algorithm with up to hundreds of thousands of layers. Our findings show that, even with reasonable computational resources per task, cluster-based approaches can significantly speed up simulations of large-sized quantum circuits and improve the fidelity of the final states.
@@ I Introduction
Quantum computing promises significant performance advantages over classical algorithms, particularly for tasks such as search [Grover1996] and factorization [Shor1999]. However, quantum computers still face serious challenges, as quantum states are vulnerable to physical effects, such as noise and decoherence, which lead to loss of information over time [NielsenChuang2010, Preskill2018NISQ]. Although significant progress has been made recently in designing efficient error-correction codes [Google2024QECbelowSCthreshold], accurately executing long-time evolutions on quantum computers remains a significant challenge.
@@
At the same time, simulating a quantum system on a classical computer poses its own challenges. In fact, a generic quantum state of a qubit system contains an amount of information that grows exponentially with the number of qubits. Consequently, the memory required to exactly represent such quantum states quickly exceeds the capabilities of any existing classical computer. This problem can be mitigated through the use of tensor network techniques [Orus2014IntroTN, Montangero2018TNReview, RevModPhys.93.045003]. In one-dimensional systems, quantum states can be efficiently represented as matrix product states (MPSs) [Fannes1992FirstMPS, Perez-Garcia2007MPS], which require significantly less memory than a full quantum state description, albeit at the cost of introducing approximation errors. By leveraging efficient MPS-based algorithms, such as time-evolving block decimation (TEBD) [Vidal2003TEBD] and the density matrix renormalization group (DMRG) [White1992DMRG, Schollwock2005DMRGRev], it is possible to achieve very low approximation errors and numerically address key problems in the study of static and dynamical properties of quantum many-body systems.
@@
The utility of tensor networks in quantum many-body physics has rapidly extended to quantum computing. While simulating ideal large-scale quantum computers on classical machines is exceedingly challenging, noisy intermediate-scale quantum (NISQ) computers exhibit imperfections that lead to computational errors comparable with the approximation errors of tensor network algorithms [Zhou2020WhatLimits]. Following Google’s claim of achieving quantum supremacy [Arute2019Sycamore] by simulating a random quantum circuit [Aaronson2017QSComplexity, Harrow2017QSupReq, Bouland2019RCS, Arute2019Sycamore] with a superconducting quantum device [Makhlin1999JJqubit, Krantz2019SuperconductingReview], a series of studies demonstrated that it was possible to simulate the same task with similar or better fidelity on large classical computers with optimized tensor network contraction techniques [Gray2021HyperExact, Pan2021simulatingsycamore, Pan2022samplingsycamore, Liu2021_304s]. Another work by the team of IBM Quantum showed evidence for the utility of NISQ devices [Preskill2018NISQ] by simulating the two-dimensional (2D) transverse-field Ising model with a Trotterized quantum circuit on a 127-qubit quantum computer [Kim2023IBMEagle]. Shortly thereafter, tensor network simulations of the same Hamiltonian were performed with good accuracy on relatively low computational resources, increasing the simulated amount of qubits up to 1121 [Tindall2024IBMTN1, Patra2024IBMTN2].
@@
Optimized classical simulations of many-body dynamics and quantum computing can be seen as complementary in the study of collective quantum effects. Indeed, optimizing numerical methods to simulate large and complex quantum systems could lead to understanding quantum phenomena at large scales, thereby aiding in the development of quantum technologies, such as creating digital twins of experimental quantum platforms [Jaschke_2024, bidzhiev2023cloudondemandemulationquantum], or optimizing TEBD and DMRG algorithms to simulate both random quantum circuits and quantum algorithms [Zhou2020WhatLimits, Ayral2023DMRG_QC, Niedermeier2024QalgsTN].
@@
In this work, we propose an approach based on entanglement clustering [Fatima2021GateCluster, Chen2021CircuitPartitioner, cicero2024simulationquantumcomputersreview] to improve tensor network algorithms for simulating quantum circuits with finite fidelity. The main idea is that, by dynamically clustering qubits, according to the pattern of entangling gates in a quantum circuit, we can optimize the simulation of deep quantum circuits with irregular entanglement distribution.
@@
In particular, we focus on the TEBD and DMRG algorithms. For the former, instead of immediately applying individual gates to the initial MPS, we perform an analysis of the structure of the simulated quantum circuit as a whole. We adaptively divide a quantum circuit into clusters of entanglement according to the amount of memory available in our classical computing system, allowing the algorithm to exactly contract multiple gate layers at the same time. We call this variant the “cluster-TEBD” algorithm. For the latter, we devise a protocol, we term the “Dynamical Adaptive Grouping Routine”, which makes use of hypergraph partitioning [Schlag2022KaHyPar] to automatically determine how to group qubits throughout the algorithm. We exploit this routine to adaptively vary the grouping in each iteration, based on the distribution of entangling gates in the contracted section of the circuit. Moreover, we employ heuristic approaches to contract multiple layers efficiently [Smith2018opt_einsum, Gray2021HyperExact].
@@
Benchmarking the original algorithms against their respective optimized variations on random-structured quantum circuits, with both Clifford and non-Clifford gates [Gottesman1998GKTheorem], as well as on a tensor network-based implementation of Shor’s algorithm for factoring [Shor1999], shows substantial improvements in both runtime and fidelity, proving the effectiveness of cluster-based optimization in finite-fidelity tensor network algorithms.
@@
The paper is organized as follows. In Sec. II, we make a comprehensive review of tensor network methods to simulate quantum circuits with finite fidelity. In Sec. III, we explain how arranging qubits in clusters of highly entangled subsystems can improve TEBD and DMRG simulations of quantum circuits. In Sec. IV, we present the results of our benchmarks, where we compare runtime and fidelity of cluster-TEBD and standard TEBD when simulating random-structured quantum circuits and Shor’s algorithm, and show the effect of contracting more layers in a single iteration of DMRG. Finally, in Sec. V, we draw our conclusions and discuss possible future directions.
@@ II Simulating quantum circuits with tensor networks
Figure 1: Quantum circuit with eight qubits and seven layers modeled as a tensor network. The initial state is an outer product of rank-1 tensors (vectors), as in this case it is a fully separable state. The gates are modeled as rank-2 tensors (matrices) for single-qubit operations, and rank-4 tensors for two-qubit operations, and they are labeled according to the qubits they act on and the layer they belong to.
In our computational framework, a tensor can be interpreted as a multidimensional array, with a number $r$ of indices, called the “rank” of the tensor, each one taking a finite number of values (“dimension”). Contracting tensors over their shared indices generates what is referred to as a “tensor network” [Orus2014IntroTN, Montangero2018TNReview]. The purpose of this section is to introduce notation and briefly illustrate how tensor networks can be used to represent and simulate quantum circuits. In particular, we present some recent developments in the context of tensor network algorithms to simulate quantum circuits [Zhou2020WhatLimits, Ayral2023DMRG_QC], introducing some quantities that will be crucial throughout the text.
@@ II.1 Mapping quantum circuits to tensor networks
NN-qubit gates in a quantum circuit can be directly represented as rank-$2N$ tensors with all indices of dimension 2. This means, for instance, that a two-qubit gate, which is usually represented as a $4\times 4$ matrix, can be mapped to a $2\times 2\times 2\times 2$ tensor. When no qubits are entangled at the start, the initial state of a circuit can also be directly represented as an outer product of rank-1 tensors (“vectors”), each of physical dimension $d=2$, corresponding to the dimension of the Hilbert space of a qubit.
@@
When dealing with high-depth quantum circuits, it is crucial to define a quantity that characterizes the position of each gate within the overall evolution. For this reason, we define a gate layer $l$ in a quantum circuit as the subsets of gates that operate on independent sets of qubits at most once at the same time step. Let us identify as $U_{(q_{1},\dots,q_{n}),l}$ a generic $n$-qubit gate applied on qubits $q_{1},\dots,q_{n}$ at layer $l$. For instance, the gates $U_{(3,4),l}$ and $U_{4,l^{\prime}}$ cannot belong to the same layer, as they operate on the same qubit $4$, making $l\neq l^{\prime}$. On the other hand, gates $U_{(3,4),l}$ and $U_{5,l^{\prime\prime}}$ could belong to the same layer, provided no operations are on qubit $5$ after $U_{(3,4),l}$. If this condition is satisfied, we can write $l=l^{\prime\prime}$. Gates belonging to the same layer can be physically interpreted as unitary operations applied at the same time step. Fig. 1 shows an example of a quantum circuit modeled as a tensor network, using the described notation for layers. Because of the way a gate layer was defined, the number of gates per layer can change depending on the structure of the circuit. In the circuit shown in Fig. 1, all layers are fully filled with gates, but if, for instance, we added two gates $U_{(6,7)}$ and $U_{(7,8)}$ at the end of the evolution, each of them would belong to two new separate layers $8$ and $9$, respectively.
[…]
II.2 Encoding entanglement: bond dimensions, truncation, and fidelity
@@
When the initial state of a quantum circuit is entangled, mapping it to tensor network notation is not as straightforward. Since the individual sites are not fully separable into tensor products, the entire state has to be stored in a tensor of size $2^{N}$, with $N$ being the number of qubits. However, such a tensor can be decomposed, as shown in Fig. 2, by performing successive singular value decompositions (SVDs) to obtain an MPS representation [Fannes1992FirstMPS, Perez-Garcia2007MPS]. Aside from the physical dimensions, this representation introduces additional virtual bond dimensions $\chi_{b}$, which encode the entanglement of the state. Assuming for simplicity $\chi_{1}=\chi_{2}=\dots=\chi$, such a representation, in principle, contains a number of parameters equal to $2N\chi^{2}$, i.e., from exponential to linear in the number of qubits. Note that a separable state can be represented equivalently as an MPS with bond dimensions $\chi_{b}=1\ \forall\ b$.
[…]
@@
In general, unitary evolutions, as those performed in quantum circuits, can lead to volume-law scaling, where the entanglement entropy scales linearly with the system size, e.g., for a one-dimensional system of size $L$:
@@
In this case, the bond dimension scales exponentially with the system size. Inevitably, the longer the evolution, the larger the bond dimension, making exact simulations extremely challenging. An approximation can be introduced by reducing the bond dimension via the truncation of singular values in the SVD step, as in Fig. 3. The singular values to truncate at each step can be determined using two parameters, namely, the cutoff $\eta$ and the maximum bond dimension $\chi_{\mathrm{max}}$. The cutoff $\eta$ determines the minimum magnitude a singular value must have to be kept, meaning all singular values smaller than $\eta$ are discarded. The maximum bond dimension $\chi_{\mathrm{max}}$ is the limit on the number of singular values to be kept, meaning if the bond dimension is larger than $\chi_{\mathrm{max}}$, all values in excess are discarded.
@@
$\mathcal{F}(n)=\lvert\braket{\psi^{\mathrm{MPS}}_{\mathrm{exact}}(n)|\psi^{\mathrm{MPS}}_{\mathrm{trunc}}(n)}\rvert^{2}.$ (3)
However, knowing the exact representation of a highly entangled quantum state requires large bond dimensions, leading back to the original problem. A way to circumvent this problem is to approximate the fidelity using singular values, as shown in Ref. [Zhou2020WhatLimits]. Let us define the two-qubit fidelity $f_{i}$ as the ratio between the sum of singular values after and before truncation:
@@
Together with runtime, the total fidelity can be used as a benchmark to evaluate the efficiency and the accuracy of classical algorithms for simulating quantum dynamics [Zhou2020WhatLimits].
@@ II.3 Algorithms for simulating quantum circuits
The main algorithms for simulating quantum circuits, originally developed for simulating static and dynamical evolution of quantum many-body systems, exploit the approximate fidelity introduced in Eq. (5) to measure the simulation error on large/highly entangled quantum states [Zhou2020WhatLimits, Ayral2023DMRG_QC].
@@
The TEBD algorithm iterates over each gate in the evolution and applies it to the corresponding qubits. Each iteration of the TEBD, analyzing a generic gate $U^{[i,i+1,\dots]}_{l+1},$ can be divided into three main steps:
@@
[…]
@@
A different approach that requires the contraction of multiple layers of the circuit is the DMRG algorithm for quantum circuit simulations, introduced by Ayral et al. [Ayral2023DMRG_QC]. The algorithm starts by defining a number of layers $L_{\mathrm{max}}$ to optimize in each step at the same time. Then, the initial MPS with tensors $A^{[1]}_{0},\dots,A^{[N]}_{0}$ is mapped to a grouped MPS with tensors $\Psi^{[1]}_{0},\dots,\Psi^{[g]}_{0}$, with $g\leq N$ the number of grouped sites. In a grouped MPS, each site $\tau$ has multiple physical indices $d_{\tau}$, depending on the number of qubits mapped to that site. Each grouped site is a tensor that represents a higher-dimensional system with a Hilbert space of dimension $2^{d_{\tau}}$, preserving all the entanglement within this group exactly without any truncation. A grouped site can be formed by contracting the sites from the original MPS that are to be combined, according to a predefined grouping scheme. The grouping scheme is predetermined prior to the algorithm’s execution and kept fixed throughout the simulation, and it is based on considerations about the structure of the circuits to simulate and on the amount of available memory. For instance, qubits on which many entangling operations are applied should belong to the same group, but the size of each grouped site should not exceed a certain threshold, as the system would run out of memory. In each step $t$, the algorithm generates a random grouped MPS $M^{[1]{\dagger}}_{tL_{\mathrm{max}}},\dots,M^{[g]{\dagger}}_{tL_{\mathrm{max}}}$ with bond dimensions $\tilde{\chi}_{i}$. Let $\Psi^{[1]}_{tL_{\mathrm{max}}},\dots,\Psi^{[g]}_{tL_{\mathrm{max}}}$ denote the MPS of the evolved state after applying $tL_{\mathrm{max}}$ layers of the circuit. The objective of the algorithm is to variationally minimize the distance between the random grouped MPS and the exact MPS representation of the state after $tL_{\mathrm{max}}$ layers. For illustrative purposes, we describe the iteration of the algorithm at step $1$ in the following:
@@
[…]
This iteration can be generalized for any step $t$, where the evolution from layer $(t-1)L_{\mathrm{max}}+1$ to layer $tL_{\mathrm{max}}$ is optimized (for further details, see Ref. [Ayral2023DMRG_QC]).